OpenStudy (abbycross167):

Can someone please help me with this question? Please? A line with slope 1/3 passes through the point (3, 6). Which point also lies on this line? a. (6, 3) b. (7, 6) c. (-3, -3) d. (-6, 3)

OpenStudy (abbycross167):

@hartnn Could you please try to help me?

hartnn (hartnn):

Point slope form of the equation of line \(y-y_1 = m (x-x_1)\) plug in m = 1/3 x1, y1 = 3,6 and get the equation first! :)

OpenStudy (abbycross167):

so y-6=1/3(x-3)?

hartnn (hartnn):

yes! now plug in each option in this, and see for which option do you get left side = right side.

hartnn (hartnn):

for ex: x,y = 6,3 3-6 = 1/3(6-3) -3 = 1/3(3) -3 = 1 which is FALSE so 6,3 is not the answer. try others

OpenStudy (abbycross167):

x,y= 7,6 3-6=1/3(7-6) -3=1/3(1) -3=1 which is FALSE so 7,6 is not the answer. Is this right?

hartnn (hartnn):

y-6=1/3(x-3) x,y = 7,6 gives 6-6 = 1/3 (7-3) continue....

OpenStudy (abbycross167):

0=1/3(4)?

hartnn (hartnn):

and thats false. so try others :)

OpenStudy (abbycross167):

@hartnn I am so sorry, my computer restarted and messed up

hartnn (hartnn):

okk, no problem...did you try 3rd option?

OpenStudy (abbycross167):

x,y= (-3, -3) -3-6=1/3(-3-3) -9=1/3(-6) -9=-6 which is FALSE so -3,-3 is not the answer. Is this right?

hartnn (hartnn):

-6/3 = -2 -9 = -2 is false. so the answer must be the 4th option!

OpenStudy (abbycross167):

okay!! thank you very much for the help!!

hartnn (hartnn):

welcome ^_^